An object of mass 0.5 kg is whirled at the end of string 0.8 m long. If the string makes three revolutions in 1.2 sec, find the tension in the string.

Grade:11

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We have 3 revolutions in 1.2 sec.
Hence, the time period of revolution would be given by:
T=\frac{1.2}{3}=0.4sec
Hence, the velocity of the object would be given by:
v=\frac{2\pi r}{T}=4\pi
So, the tension in the string would be given by:
T=\frac{mv^{2}}{r}=\frac{0.5\times 16\pi^{2}}{0.8}=98.7 N
Regards
Sumit